OpenStudy (anonymous):

Operations with Radical Questions: What is the side length b if the hypotenuse is 13 and the leg is 12?

OpenStudy (anonymous):

Halp

OpenStudy (anonymous):

sqrt( 13^2 - 12^2 )

OpenStudy (anonymous):

Could you explain further? The possible answers are: A. 1 B. 17.7 C. 5 D. 1

OpenStudy (anonymous):

I know how to use the Pythagorean Theorem (A^2 + B^2 = C^2), but I don't know how to get a leg's distance based off of the hypotenuse and another leg.

OpenStudy (anonymous):

Well, what's 13^2? And what's 12^2? And what's the difference between them?

OpenStudy (anonymous):

If the hypotenuse^2 = side1^2 + side2^2, then side2^2 would be hypotenuse^2 - side1^2

OpenStudy (anonymous):

13^2 = 169 12^2 = 144 Would the answer be 25?

OpenStudy (anonymous):

I know that isn't in the options. That's what I got.

OpenStudy (anonymous):

Close... 25 is the answer squared. What we're after is the square root of 25, and THAT will be your answer! :-)

OpenStudy (anonymous):

Thank you! I understand the process now.

OpenStudy (anonymous):

Square root of 25 is 5. Heheh.

OpenStudy (anonymous):

Excellent!! Glad I could help! And thank you! :-)

OpenStudy (anonymous):

No thank you, good sir. Have a fantastic day.